It takes Cathy 1.5 hours to paddle her canoe 6 miles upstream. Then she turns her canoe around and paddles 6 miles downstream in 1 hour.
What is the rate of the current?
What is Cathys paddling rate in still water?

Answer by stanbon(75887)

It takes Cathy 1.5 hours to paddle her canoe 6 miles upstream. Then she turns her canoe around and paddles 6 miles downstream in 1 hour.
What is the rate of the current?
What is Cathys paddling rate in still water?
------------------------
Upstream DATA:
time = 1.5 hrs ; distance = 6 miles ; rate = 6/1.5 = 4 mph
---------------
Downstream DATA:
time = 1 hr ; distance = 6 miles ; rate = 6 mph
--------------------------
Equations:
Let current rate be "c"
Let rate in still water be "w"
----
w + c = 6
w - c = 4
------
Add and solve for "w":
2w = 10
w = 5 mph (paddling rate in still water)
----
Since w+c = 6, c = 1 mph (rate of the current)
===================================================
Cheers,
stan H.